# Spectral Convolutional Conditional Neural Processes

### v3

@@ 1 Introduction
Motivated by the success of deep neural networks in large-scale function approximation, neural network-based alternatives have emerged. Neural processes (NPs; garnelo2018conditional; garnelo2018neural) exemplify this paradigm, combining ideas from GPs and deep learning within a meta-learning framework. By exposing the model to multiple realizations of an underlying stochastic process, each treated as a distinct task, NPs learn shared structure across tasks to parameterize a neural mapping that directly approximates the corresponding posterior predictive distribution (bruinsma2024convolutional). Once trained, the model enables efficient probabilistic predictions on new tasks without further optimization.
@@ 4.1 Synthetic 1-D Regression
For the sawtooth-wave benchmark, we were unable to successfully train either TNP or TETNP. Across the configurations we attempted—including increased model capacity and multiple randomized initializations—their predictions consistently collapsed to zero. We hypothesize that this failure mode stems from spectral bias, the tendency of neural networks to favor low-frequency structure over high-frequency components (rahaman2019spectral; ronen2019convergence; basri2020frequency; tancik2020fourier; fridovich2022spectral). Recent findings by vasudeva2025transformers further suggest that this bias is exacerbated in transformer architectures compared with convolutional networks. This interpretation aligns with observations by nguyen2022transformer, who originally introduced TNPs. They report degraded performance on GP samples drawn from a periodic kernel—reflected in poor log-likelihood—despite strong results on Matérn-kernel tasks. Interestingly, our experiments on periodic-kernel GP tasks (Table 1) do not replicate this limitation. A key distinction is that we employ the efficient-query TNP (EQTNP) architecture of feng2022latent, rather than the original TNP design (nguyen2022transformer) (see Appendix B.1.1 for architectural details). Finally, although sawtooth and square-wave functions share similar discontinuity and high-frequency characteristics, Table 1 shows that TNP and TETNP do not collapse when trained on square-wave signals. Investigating the source of this discrepancy falls beyond the scope of this work but represents an intriguing direction for future study.
@@ 5.1 Neural PDE Solvers
The high computational cost of traditional numerical PDE solvers has motivated the development of more efficient deep-learning–based alternatives gupta2022towards. Among these, neural operators (li2020neural; kovachki2023neural), and in particular Fourier neural operators (FNOs, li2020fourier), have become prominent. Numerous extensions have since been proposed: helwig2023group introduced rotation- and reflection-equivariant FNOs; gupta2021multiwavelet developed multiwavelet neural operators by projecting kernels onto predefined polynomial bases; and tran2021factorized reduced model complexity through separable Fourier representations. xiao2024amortized proposed the Amortized FNO (AM-FNO), which substantially cuts parameter count by amortizing the Fourier kernel parameterization, while qin2024toward analyzed FNOs through the lens of spectral bias. koshizuka2024understanding provided a mean-field theoretical perspective, and zheng2024alias introduced the Mamba Neural Operator, which couples global context via a state-space model to achieve linear complexity and representation equivalence. bartolucci2023representation examined continuous–discrete consistency, showing that it holds for convolutional neural operators (raonic2023convolutional). Additional works include orthogonal-attention eigenfunction methods for operator learning (xiao2023improved) and hierarchical transformers with frequency-aware priors for resolution-invariant super-resolution (luo2024hierarchical).
@@ 5.2 Function Space Inference
In non-parametric Bayesian modeling, GPs and deep GPs (damianou2013deep) provide flexible function-space priors with well-calibrated uncertainty estimates, but their computational cost becomes prohibitive on large datasets. This challenge has motivated the development of Bayesian neural networks (BNNs; mackay1992practical; hinton1993keeping; neal2012bayesian), which combine neural network expressiveness with Bayesian uncertainty quantification. However, specifying meaningful and scalable priors over network weights remains notoriously difficult. Recent work therefore reframes Bayesian inference in neural networks as learning a posterior over the functions induced by random weights (wolpert1993bayesian; qiu2023should). Variational implicit processes (VIPs; ma2019variational; santana2021function; ortega2022deep) generalize GPs by defining implicit stochastic processes through latent variables, while functional variational BNNs (fBNNs; sun2019functional) enforce alignment between BNN-induced priors and target priors by minimizing a functional KL divergence—though this objective can be difficult to compute and, in some cases, ill-posed (burt2020understanding). Subsequent work has aimed to address these limitations (ma2021functional; rudner2022tractable; wild2022generalized; rudner2023function; wu2024functional). More recently, several approaches have extended diffusion models to function spaces (phillips2022spectral; dutordoir2023neural; mathieu2023geometric; franzese2023continuous; lim2023score).
@@ B.1.1 Model Architectures
This section details the architectures of all CNP variants used in our experiments. Each model outputs a Gaussian predictive distribution parameterized by a mean and a pre-softplus [dugas2000incorporating] scale. The Gaussian factorizes across query points—as is standard for CNPs (see Section 2.2)—and, for multidimensional outputs, across output dimensions as well. Thus, predictions assume independence both across query locations and across components of each output vector. Unless otherwise noted, all nonlinearities use ReLU activations [nair2010rectified].
@@
In the original TNP formulation [nguyen2022transformer], all context and query embeddings are jointly processed as a single sequence by a transformer encoder, with query tokens masked to prevent them from attending to other queries or context information. This yields a quadratic complexity in the total sequence length, $\mathcal{O}\!\left((|\mathcal{D}_{c}|+|\mathcal{D}_{q}|)^{2}\right)$. Because queries never attend to one another, feng2022latent proposed a more efficient variant—the Efficient Query TNP—which we adopt here. In this design, (i) context tokens first undergo self-attention to produce updated context embeddings, and (ii) each query token is processed by cross-attending only to these updated context embeddings. This introduces a two-branch structure (contexts processed twice, queries once) but reduces the overall complexity to
$\mathcal{O}\!\left(|\mathcal{D}_{c}|^{2}+|\mathcal{D}_{q}|\,|\mathcal{D}_{c}|\right).$
Our transformer consists of six layers of multi-head attention (8 heads, head dimension 32), each followed by a feedforward subnetwork of width 256. We employ pre-norm residual blocks with layer normalization applied before both the attention and feedforward modules. All attention layers operate in a fixed input/output embedding dimension of 256. During each forward pass, every layer first performs context self-attention and then query–context cross-attention. The final-layer query embeddings are passed to a decoder MLP with two hidden layers (width 256), whose output parameterizes a Gaussian predictive distribution via $2d_{{\mathbb{Y}}}$ units corresponding to the mean and log-scale.
@@
Each grid location is processed by an MLP with two hidden layers of width 128. The resulting features are passed to a CNN based on a U-Net architecture [ronneberger2015u], consisting of six residual convolutional blocks (kernel size 11, stride 2, 128 channels) in the encoder, followed by a symmetric sequence of six transposed convolutional blocks in the decoder, with skip connections following the design of bruinsma2023autoregressive.
@@ B.2.1 Model Architectures
The model architectures and parameter counts closely follow those described in B.1.1. The only modification concerns the discretization of the functional embedding in ConvCNP and SConvCNP. Specifically, we first expand the range defined by the minima and maxima of all context and query coordinates by a margin of 0.5 in each dimension. The resulting interval is then uniformly discretized at a resolution of 48 grid points per unit length.
@@ B.2.2 Training and Data Description
We simulate trajectories using the parameter distributions in Table 8 over a total of 110 years, discarding the first 10 years as burn-in to reduce sensitivity to initial conditions. Integration is performed using the Euler–Maruyama method with time step $\Delta t=0.022$, producing 5000 solver steps over the interval $t\in[-10,100]$. Although the solver runs at this finer resolution, we record states only on a uniform grid with spacing $0.05$, yielding approximately 2200 selected time points.
@@ B.4.1 Model Architectures
As with the other models, all MLP hidden sizes are increased from $256$ to $512$. We adopt the Efficient Query TNP design described in Section B.1.1, which employs a two-branch structure in which context embeddings are processed twice and query embeddings once. In contrast to previous sections, we allocate separate parameters for the two branches: at each layer, the multi-head self-attention over contexts and the multi-head cross-attention from queries to contexts no longer share parameters. For each multi-head attention module, the feedforward width is increased from $256$ to $512$, and all input/output embedding dimensions are set to $512$. All other configurations follow Section B.1.1.
@@
Because the data lie on a regular grid, we use the on-the-grid implementation of gordon2019convolutional, eliminating the discretization step and improving computational efficiency. Let $\mathrm{I}$ denote the incomplete image, with unobserved pixels filled with dummy values, and let $\mathrm{M}_{c}$ be the binary mask indicating observed (context) pixels. For multi-channel images, the mask is broadcast along the channel dimension.
As in Section B.1.1, the convolutional deep set module produces two outputs: (i) a density channel capturing the spatial distribution of context pixels, and (ii) a kernel-smoothed representation. The kernel is implemented via a 2D convolutional layer (kernel size 11, $d_{{\mathbb{Y}}}$ input channels, 128 output channels, no bias) with a nonnegativity constraint enforced by taking absolute values of the learned weights during the forward pass [gordon2019convolutional].
The density channel is obtained by convolving this nonnegative filter with the mask $\mathrm{M}_{c}$. The kernel-smoothed component is computed by first multiplying $\mathrm{I}$ elementwise with $\mathrm{M}_{c}$ (setting non-context pixels to zero) and then applying the same nonnegative convolution. Following gordon2019convolutional, we omit normalization by the density channel, as it did not provide empirical benefits and occasionally introduced instability; however, we retain the positivity constraint on the kernel.
@@ B.4.2 Data Description
For our image-completion experiments, we use the Describable Textures Dataset (DTD; cimpoi2014describing). We adopt the standard DTD split, which contains 1,880 images each for training, validation, and testing. All images are RGB, with heights ranging from 231 to 778 pixels and widths from 271 to 900 pixels. During training, validation, and testing, we sample from each image a random $192\times 192$ crop, which we then resize to $64\times 64$. Pixel coordinates of the resulting $64\times 64$ grid are linearly mapped to [1,1][-1,1], and pixel intensities are normalized to $[0,1]$ independently across channels.
@@ B.5 Ablation Studies
We conducted ablation studies to assess the impact of three key design choices in SConvCNP:
  1. 1.
    the number of retained Fourier modes,
  2. 2.
    the discretization resolution of the functional embedding,
  3. 3.
    the use of positional encodings.
Experiments were performed on two 1D function families: samples from a Gaussian process with a Matérn–5/2 kernel and a sawtooth wave. We used a slightly smaller SConvCNP model (3.1M parameters with 32 Fourier modes), consistent with the synthetic benchmark setup. Training followed the same protocol, and evaluation was carried out on 8096 test tasks. Our findings are summarized below.

### v4

@@ 1 Introduction
Motivated by the success of deep neural networks in large-scale function approximation, neural network-based alternatives have emerged. Neural processes (NPs; garnelo2018conditional; garnelo2018neural) exemplify this paradigm, combining ideas from GPs and deep learning within a meta-learning framework. By exposing the model to multiple realizations of an underlying stochastic process, each treated as a distinct task, NPs learn shared structure across tasks to parameterize a neural mapping that directly gives an approximation of the corresponding posterior predictive distribution (bruinsma2024convolutional). Once trained, the model enables efficient probabilistic predictions on new tasks without further training.
@@ 4.1 Synthetic 1-D Regression
For the sawtooth-wave benchmark, we were unable to successfully train either TNP or TETNP. Across the configurations we attempted—including increased model capacity and multiple randomized initializations—their predictions consistently collapsed to zero. We hypothesize that this failure mode stems from spectral bias, the tendency of neural networks to favor low-frequency structure over high-frequency components (rahaman2019spectral; ronen2019convergence; basri2020frequency; tancik2020fourier; fridovich2022spectral). Recent findings by vasudeva2025transformers further suggest that this bias is exacerbated in transformer architectures compared with convolutional networks. This interpretation aligns with observations by nguyen2022transformer, who originally introduced TNPs. They report degraded performance on GP samples drawn from a periodic kernel—reflected in poor log-likelihood scores—despite strong results on Matérn-kernel tasks. Interestingly, our experiments on periodic-kernel GP tasks (Table 1) do not replicate this limitation. A key distinction is that we employ the Efficient Query TNP (EQTNP) architecture of feng2022latent, rather than the original TNP design (nguyen2022transformer) (see Appendix B.1.1 for architectural details). Finally, although sawtooth and square-wave functions share similar discontinuity and high-frequency characteristics, Table 1 shows that TNP and TETNP do not collapse when trained on square-wave signals. Investigating the source of this discrepancy falls beyond the scope of this work but represents an intriguing direction for future study.
@@ 5.1 Neural PDE Solvers
The high computational cost of traditional numerical PDE solvers has motivated the development of more efficient alternatives based on deep learninggupta2022towards. Among these, neural operators (li2020neural; kovachki2023neural), and in particular Fourier neural operators (FNOs, li2020fourier), have become popular. Numerous extensions have since been proposed: helwig2023group introduced rotation- and reflection-equivariant FNOs; gupta2021multiwavelet developed multiwavelet neural operators by projecting kernels onto predefined polynomial bases; and tran2021factorized reduced model complexity through separable Fourier representations. xiao2024amortized proposed the Amortized FNO (AM-FNO), which substantially cuts parameter count by amortizing the Fourier kernel parameterization, while qin2024toward analyzed FNOs through the lens of spectral bias. koshizuka2024understanding provided a mean-field theoretical perspective, and zheng2024alias introduced the Mamba Neural Operator, which couples global context via a state-space model to achieve linear complexity and representation equivalence. bartolucci2023representation examined continuous–discrete consistency, showing that it holds for convolutional neural operators (raonic2023convolutional). Additional works include orthogonal-attention eigenfunction methods for operator learning (xiao2023improved) and hierarchical transformers with frequency-aware priors for resolution-invariant super-resolution (luo2024hierarchical).
@@ 5.2 Function Space Inference
In non-parametric Bayesian modeling, GPs and deep GPs (damianou2013deep) provide flexible function-space priors with well-calibrated uncertainty estimates, but their computational cost becomes prohibitive on large datasets. This challenge has motivated the development of Bayesian neural networks (BNNs; mackay1992practical; hinton1993keeping; neal2012bayesian), which combine neural network expressiveness with Bayesian uncertainty quantification. However, specifying meaningful priors over network weights remains notoriously difficult. Recent work therefore reframes Bayesian inference in neural networks as learning a posterior over the functions induced by stochastic weights (wolpert1993bayesian; qiu2023should). Variational implicit processes (VIPs; ma2019variational; santana2021function; ortega2022deep) generalize GPs by defining implicit stochastic processes through latent variables, while functional variational BNNs (fBNNs; sun2019functional) enforce alignment between BNN-induced priors and target priors by minimizing a functional KL divergence—though this objective can be difficult to compute and, in some cases, ill-posed (burt2020understanding). Subsequent work has aimed to address these limitations (ma2021functional; rudner2022tractable; wild2022generalized; rudner2023function; wu2024functional). More recently, several approaches have extended diffusion models to function spaces (phillips2022spectral; dutordoir2023neural; mathieu2023geometric; franzese2023continuous; lim2023score).
@@ B.1.1 Model Architectures
This section details the architectures of all CNP variants used in our experiments. Each model outputs a Gaussian predictive distribution parameterized by a mean and a pre-softplus [dugas2000incorporating] scale. The pre-softplus value is passed through a softplus transformation, and a minimum noise term of $10^{-6}$ is then added to the resulting scale. The Gaussian factorizes across query points—as is standard for CNPs (see Section 2.2)—and, for multidimensional outputs, across output dimensions as well. Thus, predictions assume independence both across query locations and across components of each output vector. Unless otherwise noted, all nonlinearities use ReLU activations [nair2010rectified].
@@
[…]
$\mathcal{O}((|\mathcal{D}_{c}|+|\mathcal{D}_{q}|)^{2}).$
Because queries never attend to one another, feng2022latent introduced the Efficient Query TNP, which we adopt here. This variant first applies self-attention to context tokens to produce updated context embeddings, then processes each query using cross-attention over these updated contexts. The resulting two-branch structure (contexts processed twice, queries once) reduces the overall complexity to
$\mathcal{O}\bigl(|\mathcal{D}_{c}|^{2}+|\mathcal{D}_{q}|\,|\mathcal{D}_{c}|\bigr).$
Our model uses six transformer layers with eight attention heads (head dimension 32), each followed by a feedforward subnetwork of width 256. We use the standard pre-norm architecture, applying layer normalization before both the attention and feedforward modules. All attention operations share a 256-dimensional embedding dimension. At each layer, we first apply context self-attention and then query–context cross-attention. Notably, the same multi-head attention parameters are used for both operations; the distinction arises only from which embeddings serve as queries, keys, and values (yielding self-attention when they coincide, and cross-attention otherwise). This is not a modeling requirement—separate parameter sets could be used, as demonstrated in Appendix B.4.1. The final-layer query embeddings are passed to a decoder MLP with two hidden layers (width 256), whose output parameterizes a Gaussian predictive distribution via $2d_{{\mathbb{Y}}}$ units corresponding to the mean and log-scale.
@@
Each grid location is processed by an MLP with two hidden layers of width 128. The resulting features are passed to a CNN based on a U-Net architecture [ronneberger2015u], consisting of six residual convolutional blocks (kernel size 11, stride 2, 128 channels) in the encoder, followed by a symmetric sequence of six transposed convolutional blocks in the decoder, with skip connections following the design of bruinsma2023autoregressive. Because the U-Net downsamples the spatial resolution by a factor of 64, we ensure that the constructed grid size is divisible by 64. For the one-dimensional benchmarks considered here, we therefore enlarge the discretization interval when necessary so that the CNN’s output aligns exactly with the input grid size.
@@ B.2.1 Model Architectures
The model architectures and parameter counts closely follow those described in B.1.1. The only modification concerns the discretization of the functional embedding in ConvCNP and SConvCNP. Specifically, we first expand the range defined by the minima and maxima of all context and query coordinates by a margin of 0.5 in each dimension. The resulting interval is then uniformly discretized at a resolution of 48 points per unit length.
@@ B.2.2 Training and Data Description
We simulate trajectories using the parameter distributions in Table 8 over a total of 110 years, discarding the first 10 years as burn-in to reduce sensitivity to initial conditions. Integration is performed using the Euler–Maruyama method with time step $\Delta t=0.022$, producing 5000 solver steps over the interval $t\in[-10,100)$. Although the solver runs at this finer resolution, we record states only on a uniform grid with spacing $0.05$, yielding approximately 2200 selected time points spanning [10,100)[-10,100).
@@ B.4.1 Model Architectures
As with the other models, all MLP hidden sizes are increased from $256$ to $512$. We adopt the Efficient Query TNP design described in Section B.1.1, which employs a two-branch structure in which context embeddings are processed twice and query embeddings once. In contrast to previous sections, we allocate separate parameters for the two branches: at each layer, the multi-head self-attention over contexts and the multi-head cross-attention between queries and contexts no longer share parameters. For each multi-head attention module, the feedforward width is increased from $256$ to $512$, and all input/output embedding dimensions are set to $512$. All other configurations follow Section B.1.1.
@@
Because the data lie on a regular grid, we use the on-the-grid implementation of ConvCNP [gordon2019convolutional], eliminating the discretization step and improving computational efficiency. Let $\mathrm{I}$ denote an incomplete image, with unobserved pixels filled with dummy values, and let $\mathrm{M}_{c}$ be the binary mask indicating observed (context) pixels. For multi-channel images, the mask is broadcast along the channel dimension.
As in Section B.1.1, the output of the convolutional deep set module has two components: (i) a density channel capturing the spatial distribution of context pixels, and (ii) a kernel-smoothed representation. The kernel is implemented via a 2D convolutional layer (kernel size 11, $d_{{\mathbb{Y}}}$ input channels, 128 output channels, no bias) with a nonnegativity constraint enforced by taking absolute values of the learned weights during the forward pass [gordon2019convolutional].
The density channel is obtained by convolving this modified filter with the mask $\mathrm{M}_{c}$. The kernel-smoothed component is computed by first multiplying $\mathrm{I}$ elementwise with $\mathrm{M}_{c}$ (setting non-context pixels to zero) and then applying the same modified convolution. Following gordon2019convolutional, we omit normalization by the density channel, as it did not provide empirical benefits and occasionally introduced instability; however, we retain the positivity constraint on the kernel.
@@ B.4.2 Data Description
For our image-completion experiments, we use the Describable Textures Dataset (DTD; cimpoi2014describing). We adopt the standard DTD split, which contains 1,880 images each for training, validation, and testing. All images are RGB, with heights ranging from 231 to 778 pixels and widths from 271 to 900 pixels. During training, validation, and testing, we sample from each image a random $192\times 192$ crop, which we then resize to $64\times 64$. Pixel coordinates of the resulting $64\times 64$ grid along each axis are linearly mapped to [1,1][-1,1], and pixel intensities are normalized to $[0,1]$ independently across channels.
@@ B.5 Ablation Studies
We performed a series of ablation studies to quantify the contribution of three key design choices in SConvCNP: (i) the number of retained Fourier modes, (ii) the discretization resolution of the functional embedding, and (iii) the inclusion of positional encodings. Experiments were conducted on two 1D function families—samples from a Gaussian process with a Matérn–5/2 kernel and a sawtooth wave. To isolate the effect of each component, we used a slightly reduced-capacity SConvCNP model (3.1M parameters, 32 Fourier modes). Training followed the same protocol as in Section B.1, and evaluation was performed on 8,096 test tasks. A summary of our findings is provided below.